OpenStudy (anonymous):

Which shows 53^2 - 47^2 being evaluated using the difference of squares method? 53^2 - 47^2 = 2809 - 2209 = 600 53^2 - 47^2 = (53 + 47)(53 - 47) = (100)(6) = 600 53^2 - 47^2 = (53 - 47)2 = 62 = 36 53^2 - 47^2 = (2809 + 2209)(2809 - 2209) = 3,010,800

OpenStudy (kc_kennylau):

Difference of square: a^2-b^2=(a+b)(a-b)

OpenStudy (cggurumanjunath):

53^2 - 47^2 = (53 + 47)(53 - 47)

OpenStudy (kc_kennylau):

yep :)

OpenStudy (anonymous):

Thanks guys!

OpenStudy (cggurumanjunath):

wc.

OpenStudy (anonymous):

Do you guys mind helping me with one more?

OpenStudy (anonymous):

@kc_kennylau

OpenStudy (kc_kennylau):

Not at all, say it :)

OpenStudy (anonymous):

The graph of f(x) = x^2 is shifted 3 units to the right to obtain the graph of g(x). Which of the following equations best describes g(x)? g(x) = (x + 3)^2 g(x) = x^2 - 3 g(x) = x^2 + 3 g(x) = (x - 3)^2

OpenStudy (kc_kennylau):

up: y=f(x)+c down: y=f(x)-c left: y=f(x+c) right: y=f(x-c) Where y denotes the new function and c denotes how many units it is translated to.

OpenStudy (anonymous):

So what would it be? I don't get it

OpenStudy (kc_kennylau):

the fourth option :)

OpenStudy (anonymous):

ah, alright thanks! :)

OpenStudy (kc_kennylau):

no problem :)